OpenStudy (anonymous):

which equation represents a line that is perpendicular to the line represented by the equation y=3/4x-1/2 a) y=2x-3/4 b) y=3/4x+1/2…

OpenStudy (anonymous):

A line perpendicular to y=3/4x-1/2 will be of the form ... y = -4/3x +c

OpenStudy (anonymous):

Any line perpendicular to this line which has slope 3/4 must have slope -4/3 Now y = 2x - 3/4 has slope 2 and so is not perpendicular to the given line. y=3/4x+1/2 has the same slope 3/4 as the given line and is therefore parallel to the given line. So neither a) nor b) is perpendicular to the given line.